# Cost-Aware Contrastive Routing for LLMs

### v1

@@ Abstract
We study cost-aware routing for large language models across diverse and dynamic pools of models. Existing approaches often overlook prompt-specific context, rely on expensive model profiling, assume a fixed set of experts, or use inefficient trial-and-error strategies. We introduce Cost-Spectrum Contrastive Routing (CSCR), a lightweight framework that maps both prompts and models into a shared embedding space to enable fast, cost-sensitive selection. CSCR uses compact, fast-to-compute logit footprints for open-source models and perplexity fingerprints for black-box APIs. A contrastive encoder is trained to favor the cheapest accurate expert within adaptive cost bands. At inference time, routing reduces to a single $k$-NN lookup via a FAISS index, requiring no retraining when the expert pool changes and enabling microsecond latency. Across multiple benchmarks, CSCR consistently outperforms baselines, improving the accuracy–cost tradeoff by up to 25%, while generalizing robustly to unseen LLMs and out-of-distribution prompts.
@@ 1 Introduction
After a burst of reinforcement-learning and specialized finetuning, the Large Language Model (LLM)  Radford et al. (2018, 2019); Brown et al. (2020); Touvron et al. (2023); DeepSeek-AI (2024) ecosystem has fractured: code models excel at generating code but hallucinate outside programming contexts, math-tuned variants solve AIME Li et al. (2024) yet mishandle open-ended dialogue, and instruction chatbots trade being precise for fluency. Production systems therefore host a pool of models with different sizes, licenses, and domain strengths, and decide at run time which one to call for every user prompt, or worse: burden the users with picking the model they need.
A router (a.k.a. model selector, mixture-gate) adjudicates that choice online. It dynamically selects the most appropriate LLM from a pool for each input. Without it, operators either over-pay by defaulting to the largest model or risk quality regressions by pinning to cheaper ones. Current routers Hendy et al. (2023); Hari and Thomson (2023); Lu et al. (2023); Shnitzer et al. (2023) for a pool of LLMs fall into two broad camps.
Parametric routing methods, such as softmax-based “one-head-per-model” classifiers Fedus et al. (2022), optimize exclusively for top-1 accuracy without explicit consideration of inference costs. Consequently, they tend to default to selecting expensive models and require full retraining whenever new models are introduced.
Recent approaches, such as UMR Jitkrittum et al. (2025), enhance generalization by routing across a joint space of prompt clusters and model footprints. On the other hand, bandit-based methods Nguyen et al. (2024); Hu et al. (2024); Li (2025), including Thompson sampling Thompson (1933); Agrawal and Goyal (2012) and UCB Auer et al. (2002), track simplified quality–cost metrics that disregard detailed prompt characteristics, resulting in slower convergence and reduced effectiveness, especially with heterogeneous workloads. All these methods remain cost-agnostic during training, solely relying on post-hoc hyperparameter tuning to achieve an effective balance between cost and accuracy.
@@
Figure 1: Accuracy–cost/size deferral curves on three expert pools. Across all benchmarks, our Cost-Spectrum Contrastive Router (blue) consistently dominates the Pareto frontier—achieving higher accuracy at lower model size and latency (left, middle) and reduced cost (right).
@@
  • Universal ultra-compact descriptors: Two lightweight and fast to compute fingerprints that work across the full spectrum of LLMs. (1) Logit footprints that require only <10 forward passes through an open-weights model. (2) Perplexity fingerprints that score any black-box API with a small public LM, enabling vendor-agnostic routing.
  • Cost-Spectrum InfoNCE: A novel contrastive objective that (1) selects correct positives within adaptive cost bands, (2) temperature-scales each band separately, and (3) down-weights negatives in proportion to their price. This aligns the learned metric with the accuracy-cost Pareto frontier.
  • Routing Efficiency: A shared space turns routing into a single $k$-NN lookup, eliminating brittle softmax gates and retraining whenever the pool changes. Then a lightweight FAISS index makes routing to a microsecond lookup.
@@ 2.1 LLM Routing
Non-predictive methods generate outputs from one or more models before making a selection. FrugalGPT Chen et al. (2023b) uses a sequential strategy and a response quality threshold to minimize cost. Other works adopt layered inference architectures to escalate hard queries to more powerful models Wang et al. (2023), or leverage cascades with self-verification Madaan et al. (2023); Yue et al. (2023); Lee et al. (2023a); Lu et al. (2023).
@@
In contrast, predictive routing aims to select the best model before any inference is performed. Strategies include supervised learning Shnitzer et al. (2023), reward-model-based routing Hari and Thomson (2023), and meta-models trained to predict LLM performance given an input Šakota et al. (2023). Router models vary widely in implementation, including neural networks Ding et al. (2024a); Šakota et al. (2024); Chen et al. (2024); Aggarwal et al. (2024), $k$-nearest neighbors Hu et al. (2024); Shnitzer et al. (2023); Stripelis et al. (2024); Lee et al. (2023a), matrix factorization Ong et al. (2024); Zhuang et al. (2024); Li (2025), and graph neural networks Feng et al. (2024). Others incorporate model-specific tokens or train across multiple domains Devvrit et al. (2024); Cai et al. (2024a).
Academic routers usually assumed the expert set is static. Recently routing with a dynamic pool of experts has been explored Jitkrittum et al. (2025); Li (2025). UMR Jitkrittum et al. (2025) clusters probes and stores coarse capability footprints but still rebuilds them offline whenever the model pool changes. LLM-Bandit Li (2025) optimizes cost but ignores prompt semantics and offers no cold-start prior for unseen experts.
@@ 2.2 Routing within MoE and Hybrid Architectures
Routing LLMs can be viewed as a coarse-grained MoE, where each expert is a full LLM. Routing is a central mechanism in MoE models Jacobs et al. (1991); Jordan and Jacobs (1993); Shazeer et al. (2017), where expert modules are dynamically activated based on input. While classical MoEs involved equally-sized sub-models, modern approaches like Switch Transformer Fedus et al. (2022) and Mixtral Jiang et al. (2024) employ sparse activation to minimize cost. Approaches like UltraFuser Ding et al. (2024b) highlight recent advances in combining model specialization and flexibility.
@@ 2.3 Model Fusion, Merging and Cascading
Fusion strategies synthesize outputs from multiple LLMs to improve output quality Ravaut et al. (2022); Jiang et al. (2023); Guha et al. (2024); Wang et al. (2024); Lu et al. (2024b). Fusion approaches often rely on unsupervised metrics Zhang* et al. (2020); Sellam et al. (2020); Yuan et al. (2021) or ensemble voting to determine the final output Lee et al. (2023b). A related but distinct technique is model merging Lu et al. (2024a), where weights from multiple pre-trained or fine-tuned models are combined, either directly via methods like weight averaging Wortsman et al. (2022), Task Arithmetic Ilharco et al. (2023), or Fisher merging Matena and Raffel (2022). In contrast, cascading invokes models sequentially (often ordered by computational cost) and halts once a satisfactory output is generated Chen et al. (2023b); Yue et al. (2023); Gupta et al. (2024).
@@ 3 Method
This section formalizes our Cost–Spectrum Contrastive Router (CSCR) and its two drop-in, model-agnostic descriptors: logit fingerprints and perplexity fingerprints. CSCR is trained once on a fixed pool of LLMs and deployed without modification on any subset of that pool. At inference time it performs a $k$-NN lookup in a FAISS Douze et al. (2024) index  [^1] We use a FAISS IndexFlatIP. https://github.com/facebookresearch/faiss/wiki/Faiss-indexes to return the $k$ most cost–effective experts for a prompt. Throughout, let $\mathcal{H}\!=\!\{h^{(1)},\dots,h^{(M)}\}$ denote the available LLMs, $c(h)$ their normalized cost, and $\Phi(\mathbf{x})\!\in\!\mathbb{R}^{D}$ our frozen query encoder with a trainable MLP head $g_{\theta}(.)$.
@@ 3.1 Model Fingerprints
We map every LLM to a compact, task–independent vector $\mathbf{d}_{h}\!\in\!\mathbb{R}^{D^{\prime}}$. Both descriptors are gradient-free; they can be computed off-line, cached, and shipped without IP-sensitive weights.
@@ 3.1.1 Logit–Footprint Descriptors for Transparent LLMs
Equation (2) directly samples the model’s internal predictive distribution $p_{h}(\cdot)$—the very quantity trained by maximum–likelihood objective $\sum\log p_{\theta}$ Brown et al. (2020). It therefore encodes both topical preference and generation style while remaining inexpensive (only $N{\times}T$ forward passes with greedy decoding). Thus, we use (2) as the primary descriptor whenever logits are available. See Appendix C.1 for a more detailed discussion.
3.1.2 Perplexity Fingerprints for Black-Box or API-Only LLMs
Closed-source APIs (e.g., GPT-o3 Jaech et al. (2024), Gemini-2.5 Reid et al. (2024)) expose responses but hide almost all logits. For such models we adopt a per–prompt cross-entropy fingerprint:
@@ 3.1.2 Perplexity Fingerprints for Black-Box or API-Only LLMs
where $w_{1{:}L_{x}}$ are the gold target tokens (ground-truth answers if available, or probe continuations); $\operatorname{normalize}(\cdot)$ denotes mean-centering and unit–variance scaling. Practically, we approximate (3) with a lightweight open LM that scores the API output $\hat{y}_{h}(x)$ instead of inaccessible $p_{h}$:
@@
Cross-entropy is proportional to KL($p_{\text{true}}\!\parallel\,p_{h}$) plus entropy of the data distribution; it therefore quantifies text fit and has long been a proxy for LM quality Mikolov and Zweig (2012); Radford et al. (2019). Prior work has shown that perplexity can serve as an effective metric for distinguishing between human-generated text and LLM-generated output Hans et al. (2024). Moreover, although $\tilde{\ell}_{h}$ is an approximation, the descriptor vector in (4) still captures how hard each prompt is for a given expert; routing on these vectors recovers much of the benefit of logit footprints while remaining viable for black-box LLMs. Figure 2 shows that, despite using the same model to compute the perplexity of text generated by different LLMs, we still observe a clear separation between the expert descriptors.
@@
We can make both descriptors reside on the unit sphere $\mathbb{S}^{K-1}$ by setting $N=K$ (i.e., the number of top tokens equals the number of prompts) with cosine similarity $\sigma(\mathbf{d}_{1},\mathbf{d}_{2})=\mathbf{d}_{1}^{\top}\mathbf{d}_{2}$. A single $k$-NN router can thus mix transparent experts (logit fingerprints) and opaque experts (perplexity fingerprints) without altering downstream training loss; only the descriptor extraction pipeline changes (See Table 4) .
3.2 Cost-Spectrum Contrastive Router
A contrastive router learns a shared embedding space where each query vector is pulled toward the descriptor of the right-sized expert and pushed away from less suitable ones; this yields three key pay-offs. First, because routing reduces to a nearest-neighbor lookup in that space, inference is a microsecond operation that adds virtually no latency to serving large expert pools. Second, contrastive objectives let the router exploit implicit supervision (“expert X solved this prompt while expert Y failed” or “X is cheaper than an equally accurate Y”) so it can be trained with only correctness or cost signals, no dense human annotations. Third, the geometry learned by contrastive learning naturally generalizes: queries that look semantically or structurally similar land near the same expert regions, which improves robustness to distribution shift and unseen prompts, a behavior long noted in contrastive representation learning. Together, these properties make a contrastive router an efficient, supervision-light and highly adaptable choice for directing traffic in modern multi-LLM systems.
@@ 3.2 Cost-Spectrum Contrastive Router
Let a minibatch contain $B$ queries $\{\mathbf{x}_{i}\}_{i=1}^{B}$ and a memory bank of $M$ keys $\{\mathbf{e}_{m}\}_{m=1}^{M}$. A query encoder $f_{\theta}\!:\!\mathcal{X}\!\rightarrow\!\mathbb{R}^{d}$ produces representations $\mathbf{q}_{i}=f_{\theta}(\mathbf{x}_{i})/\hskip 2.0pt\!\|\!f_{\theta}(\mathbf{x}_{i})\!\|_{2}$, and the keys are $\ell_{2}$-normalized in advance, $\mathbf{e}_{m}=\mathbf{E}_{m}/\|\mathbf{E}_{m}\|_{2}$ (here, $E_{m}$ is the expert descriptor from Equation (2) or  (4), i.e. $E_{m}=d(h_{m})$). For each query $i$ let $\mathcal{P}(i)\subset[M]$ denote the positives (e.g. correct experts) and $\mathcal{N}(i)=[M]\setminus\mathcal{P}(i)$ the negatives. The vanilla InfoNCE objective van den Oord et al. (2019) maximizes a log-softmax over cosine similarities
@@
3.4 Cost-Spectrum InfoNCE.
@@
We therefore introduce a cost-aware spectrum version that:
@@
    Assigns band-specific temperatures so that harder (costlier) positives yield smoother gradients
@@
Cost-dependent temperature.
@@
Thus the contrastive potential aligns three signals in the same metric space: (i) semantic proximity via the query encoder, (ii) expert capability via fingerprints, and (iii) user preference via cost scaling. Previous cost-aware objectives for retrieval weight the final scoring function at inference time (e.g. Li (2025); Jitkrittum et al. (2025)); our formulation also integrates cost during representation learning, inducing a feature geometry that naturally interpolates accuracy and cost. Equation (8) collapses to standard InfoNCE when K=1K\!=\!1 and $\lambda\!=\!0$. See Appendices  C.3 and C.4 for a more detailed discussion.

### v2

@@ Abstract
We study cost-aware routing for large language models across diverse and dynamic pools of models. Existing approaches often overlook prompt-specific context, rely on expensive model profiling, assume a fixed set of experts, or use inefficient trial-and-error strategies. We introduce Cost-Spectrum Contrastive Routing (CSCR), a lightweight framework that maps both prompts and models into a shared embedding space to enable fast, cost-sensitive selection. CSCR uses compact, fast-to-compute logit footprints for open-source models and perplexity fingerprints for black-box APIs. A contrastive encoder is trained to favor the cheapest accurate expert within adaptive cost bands. At inference time, routing reduces to a single kk‑NN lookup via a FAISS index, requiring no retraining when the expert pool changes and enabling microsecond latency. Across multiple benchmarks, CSCR consistently outperforms baselines, improving the accuracy–cost tradeoff by up to 25%, while generalizing robustly to unseen LLMs and out-of-distribution prompts.
@@ 1 Introduction
After a burst of reinforcement‑learning and specialized finetuning, the Large Language Model (LLM)  [65, 66, 6, 86, 17] ecosystem has fractured: code models excel at generating code but hallucinate outside programming contexts, math‑tuned variants solve AIME [46] yet mishandle open‑ended dialogue, and instruction chatbots trade being precise for fluency. Production systems therefore host a pool of models with different sizes, licenses, and domain strengths, and decide at run time which one to call for every user prompt, or worse: burden the users with picking the model they need.
A router (a.k.a. model selector, mixture‑gate) adjudicates that choice online. It dynamically selects the most appropriate LLM from a pool for each input. Without it, operators either over‑pay by defaulting to the largest model or risk quality regressions by pinning to cheaper ones. Current routers [33, 32, 51, 76] for a pool of LLMs fall into two broad camps.
Parametric routing methods, such as softmax-based “one-head-per-model” classifiers [24], optimize exclusively for top-1 accuracy without explicit consideration of inference costs. Consequently, they tend to default to selecting expensive models and require full retraining whenever new models are introduced.
Recent approaches, such as UMR [40], enhance generalization by routing across a joint space of prompt clusters and model footprints. On the other hand, bandit-based methods [61, 34, 48], including Thompson sampling [85, 2] and UCB [3], track simplified quality–cost metrics that disregard detailed prompt characteristics, resulting in slower convergence and reduced effectiveness, especially with heterogeneous workloads. All these methods remain cost-agnostic during training, solely relying on post-hoc hyperparameter tuning to achieve an effective balance between cost and accuracy.
@@
Figure 1: Accuracy–cost/size deferral curves on three expert pools. Across all benchmarks, our Cost‑Spectrum Contrastive Router (blue) consistently dominates the Pareto frontier—achieving higher accuracy at lower model size and latency (left, middle) and reduced cost (right).
@@
  • Universal ultra-compact descriptors: Two lightweight and fast to compute fingerprints that work across the full spectrum of LLMs. (1) Logit footprints that require only <10 forward passes through an open‑weights model. (2) Perplexity fingerprints that score any black‑box API with a small public LM, enabling vendor‑agnostic routing.
  • Cost-Spectrum InfoNCE: A novel contrastive objective that (1) selects correct positives within adaptive cost bands, (2) temperature‑scales each band separately, and (3) down‑weights negatives in proportion to their price. This aligns the learned metric with the accuracy‑cost Pareto frontier.
  • Routing Efficiency: A shared space turns routing into a single kk‑NN lookup, eliminating brittle softmax gates and retraining whenever the pool changes. Then a lightweight FAISS index makes routing to a microsecond lookup.
@@ 2.1 LLM Routing
Non-predictive methods generate outputs from one or more models before making a selection. FrugalGPT [10] uses a sequential strategy and a response quality threshold to minimize cost. Other works adopt layered inference architectures to escalate hard queries to more powerful models [91], or leverage cascades with self-verification [54, 98, 43, 51].
@@
In contrast, predictive routing aims to select the best model before any inference is performed. Strategies include supervised learning [76], reward-model-based routing [32], and meta-models trained to predict LLM performance given an input [71]. Router models vary widely in implementation, including neural networks [19, 88, 11, 1], $k$-nearest neighbors [34, 76, 80, 43], matrix factorization [62, 105, 48], and graph neural networks [25]. Others incorporate model-specific tokens or train across multiple domains [18, 7].
Academic routers usually assumed the expert set is static. Recently routing with a dynamic pool of experts has been explored [40, 48]. UMR [40] clusters probes and stores coarse capability footprints but still rebuilds them offline whenever the model pool changes. LLM‑Bandit [48] optimizes cost but ignores prompt semantics and offers no cold-start prior for unseen experts.
@@ 2.2 Routing within MoE and Hybrid Architectures
Routing LLMs can be viewed as a coarse-grained MoE, where each expert is a full LLM. Routing is a central mechanism in MoE models [36, 41, 75], where expert modules are dynamically activated based on input. While classical MoEs involved equally-sized sub-models, modern approaches like Switch Transformer [24] and Mixtral [38] employ sparse activation to minimize cost. Approaches like UltraFuser [20] highlight recent advances in combining model specialization and flexibility.
@@ 2.3 Model Fusion, Merging and Cascading
Fusion strategies synthesize outputs from multiple LLMs to improve output quality [68, 39, 29, 89, 52]. Fusion approaches often rely on unsupervised metrics [99, 73, 97] or ensemble voting to determine the final output [44]. A related but distinct technique is model merging [50], where weights from multiple pre-trained or fine-tuned models are combined, either directly via methods like weight averaging [92], Task Arithmetic [35], or Fisher merging [56]. In contrast, cascading invokes models sequentially (often ordered by computational cost) and halts once a satisfactory output is generated [10, 98, 30].
@@ 3 Method
This section formalizes our Cost–Spectrum Contrastive Router (CSCR) and its two drop‑in, model‑agnostic descriptors: logit fingerprints and perplexity fingerprints. CSCR is trained once on a fixed pool of LLMs and deployed without modification on any subset of that pool. At inference time it performs a $k$-NN lookup in a FAISS [21] index  [^1] We use a FAISS IndexFlatIP. https://github.com/facebookresearch/faiss/wiki/Faiss-indexes to return the $k$ most cost–effective experts for a prompt. Throughout, let $\mathcal{H}\!=\!\{h^{(1)},\dots,h^{(M)}\}$ denote the available LLMs, $c(h)$ their normalized cost, and $\Phi(\mathbf{x})\!\in\!\mathbb{R}^{D}$ our frozen query encoder with a trainable MLP head $g_{\theta}(.)$.
@@ 3.1 Model Fingerprints
We map every LLM to a compact, task–independent vector $\mathbf{d}_{h}\!\in\!\mathbb{R}^{D^{\prime}}$. Both descriptors are gradient‑free; they can be computed off‑line, cached, and shipped without IP‑sensitive weights.
@@ 3.1.1 Logit–Footprint Descriptors for Transparent LLMs
Equation (2) directly samples the model’s internal predictive distribution $p_{h}(\cdot)$—the very quantity trained by maximum–likelihood objective $\sum\log p_{\theta}$ [6]. It therefore encodes both topical preference and generation style while remaining inexpensive (only $N{\times}T$ forward passes with greedy decoding). Thus, we use (2) as the primary descriptor whenever logits are available. See Appendix C.1 for a more detailed discussion.
3.1.2 Perplexity Fingerprints for Black‑Box or API‑Only LLMs
Closed‑source APIs (e.g., GPT‑o3 [37], Gemini-2.5 [69]) expose responses but hide almost all logits. For such models we adopt a per–prompt cross‑entropy fingerprint:
@@ 3.1.2 Perplexity Fingerprints for Black‑Box or API‑Only LLMs
where $w_{1{:}L_{x}}$ are the gold target tokens (ground‑truth answers if available, or probe continuations); $\operatorname{normalize}(\cdot)$ denotes mean‑centering and unit–variance scaling. Practically, we approximate (3) with a lightweight open LM that scores the API output $\hat{y}_{h}(x)$ instead of inaccessible $p_{h}$:
@@
Cross‑entropy is proportional to KL($p_{\text{true}}\!\parallel\,p_{h}$) plus entropy of the data distribution; it therefore quantifies text fit and has long been a proxy for LM quality [58, 66]. Prior work has shown that perplexity can serve as an effective metric for distinguishing between human-generated text and LLM-generated output [31]. Moreover, although $\tilde{\ell}_{h}$ is an approximation, the descriptor vector in (4) still captures how hard each prompt is for a given expert; routing on these vectors recovers much of the benefit of logit footprints while remaining viable for black‑box LLMs. Figure 2 shows that, despite using the same model to compute the perplexity of text generated by different LLMs, we still observe a clear separation between the expert descriptors.
@@
We can make both descriptors reside on the unit sphere $\mathbb{S}^{K-1}$ by setting $N=K$ (i.e., the number of top tokens equals the number of prompts) with cosine similarity $\sigma(\mathbf{d}_{1},\mathbf{d}_{2})=\mathbf{d}_{1}^{\top}\mathbf{d}_{2}$. A single kk‑NN router can thus mix transparent experts (logit fingerprints) and opaque experts (perplexity fingerprints) without altering downstream training loss; only the descriptor extraction pipeline changes (See Table 4) .
3.2 Cost‑Spectrum Contrastive Router
A contrastive router learns a shared embedding space where each query vector is pulled toward the descriptor of the right‑sized expert and pushed away from less suitable ones; this yields three key pay‑offs. First, because routing reduces to a nearest‑neighbor lookup in that space, inference is a microsecond operation that adds virtually no latency to serving large expert pools. Second, contrastive objectives let the router exploit implicit supervision (“expert X solved this prompt while expert Y failed” or “X is cheaper than an equally accurate Y”) so it can be trained with only correctness or cost signals, no dense human annotations. Third, the geometry learned by contrastive learning naturally generalizes: queries that look semantically or structurally similar land near the same expert regions, which improves robustness to distribution shift and unseen prompts, a behavior long noted in contrastive representation learning. Together, these properties make a contrastive router an efficient, supervision‑light and highly adaptable choice for directing traffic in modern multi‑LLM systems.
@@ 3.2 Cost‑Spectrum Contrastive Router
Let a minibatch contain $B$ queries $\{\mathbf{x}_{i}\}_{i=1}^{B}$ and a memory bank of $M$ keys $\{\mathbf{e}_{m}\}_{m=1}^{M}$. A query encoder $f_{\theta}\!:\!\mathcal{X}\!\rightarrow\!\mathbb{R}^{d}$ produces representations $\mathbf{q}_{i}=f_{\theta}(\mathbf{x}_{i})/\hskip 2.0pt\!\|\!f_{\theta}(\mathbf{x}_{i})\!\|_{2}$, and the keys are $\ell_{2}$‑normalized in advance, $\mathbf{e}_{m}=\mathbf{E}_{m}/\|\mathbf{E}_{m}\|_{2}$ (here, $E_{m}$ is the expert descriptor from Equation (2) or  (4), i.e. $E_{m}=d(h_{m})$). For each query $i$ let $\mathcal{P}(i)\subset[M]$ denote the positives (e.g. correct experts) and $\mathcal{N}(i)=[M]\setminus\mathcal{P}(i)$ the negatives. The vanilla InfoNCE objective [87] maximizes a log‑softmax over cosine similarities
@@
3.4 Cost‑Spectrum InfoNCE.
@@
We therefore introduce a cost‑aware spectrum version that:
@@
    Assigns band‑specific temperatures so that harder (costlier) positives yield smoother gradients
@@
Cost‑dependent temperature.
@@
Thus the contrastive potential aligns three signals in the same metric space: (i) semantic proximity via the query encoder, (ii) expert capability via fingerprints, and (iii) user preference via cost scaling. Previous cost‑aware objectives for retrieval weight the final scoring function at inference time (e.g. [48, 40]); our formulation also integrates cost during representation learning, inducing a feature geometry that naturally interpolates accuracy and cost. Equation (8) collapses to standard InfoNCE when K=1K\!=\!1 and $\lambda\!=\!0$. See Appendices  C.3 and C.4 for a more detailed discussion.